OpenStudy (anonymous):

Which related equation can be used to solve the equation? Column A 1. –8.7 = k + 6 2. –6 = k + 8.7 3. 8.7 = k + (–6) 3. 8.7 = k + (–6) 4. 6 = k + 8.7 Column B A. k = –8.7 – 6 B. k = 6 – 8.7 C. k = 8.7 – (–6)

OpenStudy (anonymous):

Hi, Welcome to open study!!!

OpenStudy (anonymous):

for each question 1-4 doing them one at a time solve for k make k the subject remember to keep the balance what you do to one side of an equation also do to the other . you have 1 -4 k = whatever now look in column be and find the choice that matches. for example 2. –6 = k + 8.7 k is on the right to get k only -8.7 keep the balance so -8.7 from the left also you get -6 -8.7 = k which matches option A solution 2 A

OpenStudy (anonymous):

Ehh just tell me the anwser this my 70th acc i made

OpenStudy (anonymous):

JUST TELL ME THE DAM ANWSER I AREADY SAW THAT POST IDGAF BOUT IT I WANT W/ GOES WHERE ACTUAL ANWSER NOT TO FUGURE IT OUT MYSELF

OpenStudy (anonymous):

ffs u ppl

OpenStudy (anonymous):

Wow watch the words!!! Ok I will give you the answer A, You can a little nicer you know!

OpenStudy (anonymous):

No I cant i have to finsh my stpd math so i can go back and play black ops 3

OpenStudy (anonymous):

@Vampirelove28 Yeah go F u ck yourself b1tch

OpenStudy (anonymous):

OK, I get all you wont to do is go play video games! but We cant just give you answers you haft to understand to!

OpenStudy (anonymous):

No i fluttering dont im home shcooled as fluttering long as i past the grade idgaf homeshcooling dosnt have state test and high shcool sht to pass Just get every thing right ffs

OpenStudy (anonymous):

Now get me the fluttering anwsers and stfu

OpenStudy (anonymous):

Why when i said f uc k it makes it fluttering...

OpenStudy (anonymous):

ima play mc while i wait for your dumb @ss to get me my anwsers

OpenStudy (anonymous):

OK, IF YOU DONT STOP WITH THE WORDS! AND CLAM DOWN I WILL REPORT YOU!!!

OpenStudy (anonymous):

FKING G00D REPORT ME THIS NOT MY FIRST GOT THAT GOT DELETED FFS THIS MY 70TH d umb @ss

OpenStudy (anonymous):

I am not going to give the answer to you!!! until you can stop!

OpenStudy (anonymous):

k go do it now (Slave)

OpenStudy (anonymous):

@YoungStudier @tyler_peters

OpenStudy (anonymous):

@Bookworm14 hELP

OpenStudy (anonymous):

I am not your slave

OpenStudy (anonymous):

@Vampirelove28 JUST flutterING SOLVE IT AND TELL ME ANWSER GOD YOUR ANNOYING

OpenStudy (anonymous):

idek

OpenStudy (anonymous):

b answers

OpenStudy (bookworm14):

I am sorry love, I do not give out direct answers. @Vampirelove28 you cannot help someone who does not want to be helped xD stick to your morals We can get kicked off the site for giving direct answers, and calm down! There is no need for all of the foul language, there are young kids on here!

OpenStudy (anonymous):

WHo give a fu ck bout the kids and If i wanted to kno how to do it think id fking read the book now give me direct anwser

OpenStudy (anonymous):

@NatsuOniKaze

OpenStudy (anonymous):

5 hours later 2 dumb A$$ trying to solve a equation

OpenStudy (bookworm14):

Sweety, I don't know who you think you are nor who you are talking to, but you need to be taught manners and respect! @ganeshie8 @Michele_Laino

OpenStudy (anonymous):

Please Stop with these words!!!

OpenStudy (anonymous):

He is talking to me @Vampirelove28 and you @Bookworm14

OpenStudy (bookworm14):

@Vampirelove28 just ignore him. I am not playing his game, so I am just ignoring :D I refuse to be as immature as this little kid lol Im to old for all this crap xD

OpenStudy (anonymous):

@TheTh1nker

OpenStudy (anonymous):

KK

OpenStudy (anonymous):

Dont disrespect @Vampirelove28, You can work it out youre self, And this is the end of it.

OpenStudy (anonymous):

Thanks @TheTh1nker